A wholesaler sells three types three types of GPS systems. A dealer orders twice as many economy systems at $40 each and 75 more econoplus systems at $80 dollar each than deluxe systems at $140 each for $42,000. How many of each were ordered?
Answer by Boreal(15235)

40E+75P+140D=42000
P=D+75
E=2D
40(2D)+80(D+75)+140D=42000
80D+80D+6000+140D=42000=300D+6000
300D=36000
D=120 deluxe ($16800)
P=195 plus ($15600)
E=240 economy ($9600)